ISS Daily Summary Report – 06/02/14

Ocular Health (OH): Commander (CDR) Swanson conducted his Flight Day (FD) 60 Ocular Health vision test. Flight Engineer (FE)-6 Gerst and Swanson took turns using the tonometer to collect eye pressure measurements. Tomorrow, retinal images will be collected with the fundoscope. The Ocular Health protocol calls for a systematic gathering of physiological data to characterize the risk of microgravity-induced visual impairment/intracranial pressure in ISS crewmembers. Researchers believe that the measurement of visual, vascular and central nervous system changes over the course of this experiment and during the subsequent post-flight recovery will assist in the development of countermeasures, clinical monitoring strategies, and clinical practice guidelines.

Advanced Colloids Experiment (ACE)-M-2 Sample Mixing: FE-5 Wiseman mixed ACE-M-2 samples in preparation for sample analysis in the Fluids Integrated Rack (FIR) at a later date.  In the ACE-M-2 investigation, a model colloidal system is used to observe the microscopic behavior of liquids and gases separating from each other, near the critical point, via a process called spinodal decomposition.

Emergency Roles & Responsibilities (R&R) Review/Crew Handover: All 6 crewmembers reviewed emergency general instructions and discussed overall emergency response. Following this activity, Swanson continued general handover with the 39S crew.

Japanese Pressurized Module (JPM) Return Grille Cleaning: Swanson removed return grille mesh covers and cleaned inside the grilles in the JPM.

Robotics Operations:  Robotics Ground Controllers powered up the Mobile Servicing System (MSS) and maneuvered the Space Station Remote Manipulator System (SSRMS) to a Mobile Transporter (MT) translation configuration.  This also moved the SSRMS out of a keepout zone for the SCAN Testbed payload.  Controllers also verified that the iris of the Camera Light Pan/Tilt Unit Assembly (CLPA) installed on SSRMS Boom B last Tuesday would close nominally.

Positive High Beta:  The ISS is in a period of high solar beta, with a peak of +74 degrees on June 4.  The ISS will be continuous sunlight until June 6.  Ground teams are closely monitoring thermal data, and no hardware has exceeded thermal limits.

Today’s Planned Tasks

  • RSS1,2 reboot / r/g 1009 [Complete]
  • СМ-У connection inspection at МНР-НС behind АСУ panel 139 [Complete]
  • Laptop RS1(2) reboot [Complete]
  • SM ПСС Test / РСУ, [Complete]
  • MO-8. Setup [Complete]
  • Body mass measurement  [Complete]
  • MO-8. Close-out ops [Complete]
  • DPC (S-band) [Complete]
  • Ocular Health (OH) – ocular health examination [Complete]
  • Stow returnable equipment into Soyuz 421 (h/w compartment) r/g 5685, 5750 [Complete]
  • Daily electrocardiography (start) / r/g 5744 [Complete]
  • Working СВ camera setup [Complete]
  • FMK atmospheric formaldehyde monitor deployment [Complete]
  • ACE. Science program review[Complete]
  • Ocular health questionnaire [Complete]
  • ACE. Sample stirring for 40 min. [Complete]
  • Ocular Health (OH) – ocular health examination [Complete]
  • NAPOR-mini РСА. [БЗУ-М] s/w upgrade. r/g 5769 [Complete]
  • SLT reboot [Complete]
  • Ocular Health (OH). Tonometer setup [Complete]
  • Ocular health questionnaire [Complete]
  • Ocular Health (OH). Tonometer – agent [Complete]
  • MOTOKARD. Experiment run. r/g 5763 [Complete]
  • Ocular Health (OH). Tonometer – subject [Complete]
  • Ocular Health (OH). Tonometer – agent [Complete]
  • Ocular Health (OH). Tonometer – subject [Complete]
  • MOTOKARD. Science ops assistance / r/g 5764 [Complete]
  • Ocular Health (OH). Tonometer stow [Complete]
  • BIOSIGNAL. Cryogem-03 temperature monitoring / r/g 5760  [Complete]
  • ACE. Sample stirring for 40 min. [Complete]
  • СОЖ maintenance  [Complete]
  • Crew adaptation and station review time [Complete]
  • БВН grid cleaning in Soyuz 712 [БО] maintenance compartment  [Complete]
  • Transfer into Progress 713 with IMS updates / r/g 5735 [Complete]
  • СВ mode adjustment [Complete]
  • Sampling into GSC container [Complete]
  • СВ xenon lamp c/o [Complete]
  • ACE. Sample stirring for 40 min. [Complete]
  • BSA battery cycling start [Complete]
  • СВ – closeout prep [Complete]
  • VEG – daily vegetation check [Complete]
  • VEG-01 – fill in water to water the vegetation [Complete]
  • Urine PH Kit Trash [Complete]
  • REMOTE RS LAPTOP setup / r/g 5757 [Complete]
  • Emergency roles and responsibilities review [Complete]
  • Adjust water separation safety valve and perform test separation into ЕДВ via [МФР]. r/g 5755 [Complete]
  • ISS Crew Handover [Complete]
  • JRNL – log update [Complete]
  • SHD – daily headache monitoring questionnaire [Complete]
  • IMS update [Complete]
  • Monthly vacuuming of БРИ h/w surface with power OFF and cable network connected. r/g 5373  [Complete]
  • Crew adaptation and station review time [Complete]
  • WRM – condensate transfer start [Complete]
  • Crew adaptation and station review time [Complete]
  • Exercise data dump / r/g 2639 [Complete]
  • On MCC Go Soyuz 712 [СтА] cover seal drive action c/o. r/g 5756 [Complete]
  • WRM – condensate transfer completion [Complete]
  • SHD – daily headache monitoring questionnaire [Complete]
  • Crew adaptation and station review time [Complete]
  • JEM protective grid cleaning [Complete]
  • HRF – h/w setup [Complete] 

Task List

  • ARED Cylinder Evac
  • WHC KTO replace
  • Roscosmos Report
  • Uragan photo
  • 3DA1-Cam-Ops
  • BASS ops overview
  • BASS single TL ops parts 1,2,3
  • BCATC1 sample 8 pre-mix
  • CEVIS Isolator MOD
  • Cygnus pre-gather
  • Crew handover
  • CFE ICF5 ops parts 1,2,3
  • EMU LLB Auto-Init
  • EMU LLB SSC Relocate
  • EMU LLB SSC Return
  • MERLIN PAS remove
  • EVA ORU Bag Inspection
  • IFM MOOT restow
  • NOD2O2 Cleanout
  • JPM HD assessment
  • No-skid tape audit
  • NOD3 PEPS assessment
  • P/TV Hardware questions
  • PMM cleanup [Complete]
  • WRS MRF remove
  • P/TV h/w questions part 2
  • Possum PL photo
  • RS EVA h/w gather


  • Nominal commanding
  • SSRMS maneuver [Complete]

Three-Day Look Ahead:

  • Tuesday, 06/03: Ocular Health, Crew handover, EMU PWR sample and drain
  • Wednesday, 06/04: ACE, Ocular Health, Cygnus HCP setup, EMU Antenna cable install
  • Thursday, 06/05: EMU 3010 Return To Service, Prox HCP checkout, LA1B RPCM R&Rs

QUICK ISS Status – Environmental Control Group:

                              Component Status
Elektron On
Vozdukh Manual
[СКВ] 1 – SM Air Conditioner System (“SKV1”) Off
[СКВ] 2 – SM Air Conditioner System (“SKV2”) Off
Carbon Dioxide Removal Assembly (CDRA) Lab Override
Carbon Dioxide Removal Assembly (CDRA) Node 3 Operate
Major Constituent Analyzer (MCA) Lab Stop
Major Constituent Analyzer (MCA) Node 3 Operate
Oxygen Generation Assembly (OGA) Standby
Urine Processing Assembly (UPA) Standby
Trace Contaminant Control System (TCCS) Lab Off
Trace Contaminant Control System (TCCS) Node 3 Full up